16. Find the orthogonal trajectories for the family of curves represented by 2x² + y2 = k, where k is a parameter.
Answer #1

2x² + y² = - o Differentiating with respect to a we get 4x + 2 y d y = 0 2x + y d y = 0 - ③ is the differential equation of f